# Beyond May: Complexity-stability relationships in disordered dynamical systems

### v1

@@ Abstract
Robert May famously used random matrix theory to predict that large, complex systems cannot admit stable fixed points. However, this general conclusion is not supported by empirical observation: from cells to biomes, biological systems are large, complex—and, by and large, stable. In this paper, we revisit May’s argument in the light of recent developments in both ecology and random matrix theory. Using a non-linear generalization of the Lotka-Volterra model, we show that there are in fact two kinds of complexity-stability relationships in disordered dynamical systems: if self-interactions grow faster with density than cross-interactions, complexity is destabilizing; but if cross-interactions grow faster than self-interactions, complexity is stabilizing. Our result shows that May’s principle that ”complexity begets instability” is not a general property of complex systems; instead, it is a property of a subclass of weakly cross-regulated disordered systems.
@@
Introduction.—Few mathematical arguments have influenced biological thinking like May’s prediction that large, complex ecosystems cannot be stable [1]. It is, on the face of it, a perplexing conclusion. On the one hand, May’s mathematical argument is simple and seemingly model-free, suggesting universal applicability, also beyond biology [2, 3]. On the other hand, it is clear that at least some large, complex systems are stable—else which regularities would biology be studying in the first place? In fact, empirical observation suggests the opposite relationship between complexity and stability: species-rich, strongly-coupled communities such as rainforests tend to be stable over time, while sparser ones, for instance arctic communities, often exhibit large fluctuations, extinctions, and invasions [4, 5, 6]. The tension between May’s theoretical argument and observation is at the center of the longstanding “diversity-stability debate” in ecology [7, 8].
@@
May’s argument can be summarized as follows [1]. Consider a system with $N$ populations $x_{i}$, characterized by an equilibrium point $\mathbf{x}^{*}$. Near that equilibrium $\mathbf{x}=\mathbf{x}^{*}+\delta\mathbf{x}$, the dynamics of the system is described by linear equations $d(\delta\mathbf{x})/dt=A(\delta\mathbf{x})$, and the stability of these equations requires that all eigenvalues of $A$ have negative real part. But if $A$ can be represented as $A=B-I$, where $B$ consists of random, independent interactions (with zero mean and variance $\sigma^{2}$), and $-I$ corresponds to stabilizing self-interactions on some natural timescale, then the circular law of random matrix theory implies that all eigenvalues of $A$ will have negative real part only if $\sigma^{2}N<1$. This places a sharp constraint on both diversity $N$ and interaction strength $\sigma$, often referred to as “complexity begets instability”. (This argument generalizes to $\langle B_{ij}\rangle\neq 0$, incomplete connectivity, or correlated interactions [9].)
@@
Many authors have sought to ease the tension between May’s prediction and empirical observation by invoking effects not captured by dynamical systems with random coefficients [7, 10, 11, 12, 13, 14]. In this letter, we use recent results in the physics of disordered systems [15, 16] to show that May’s argument itself is incomplete: in random dynamical systems, stability does not necessarily decrease with dimensionality and interaction strength—the opposite behavior is also possible, without the need for special or additional structure. For an in-depth discussion of the ecological implications of our result, see [17].
@@
Here $f(x_{i})$ represents the self-dynamics of a population $i$ (growth and self-regulation), while $g(x_{i})$ and $h(x_{j})$ capture the cross-interaction of $i$ with other populations. That interaction is weighted by a coefficient $A_{ij}$, such that $A_{ij}>0$ (resp. $A_{ji}<0$) implies a negative (resp. positive) effect of $j$ on the growth of $i$. Non-zero diagonal elements $A_{ii}$ accounts for additional self-interactions with the same functional form as cross-interactions. We refer to $f$ as the “production function” and $g$ as the “response function”.
This general setting has been used to study universality in network dynamics [18] and to construct minimal models of complex dynamics [19], with applications ranging from biological [20, 21] to social [22, 23, 24] systems. The classic generalized Lotka-Volterra (GLV) competition model studied recently by Bunin and collaborators [25, 26] corresponds to $f$, $g$, $h$ all linear.
However, it is natural both physically and biologically to consider more general functions, including power laws $f(x)\sim x^{\alpha}$, $g(x)\sim x^{\beta}$, $h(x)\sim x^{\gamma}$. From a physical perspective, we can imagine populations $x_{i}$ forming three-dimensional clusters whose growth is limited to their two-dimensional surface, leading to a production function $f(x)\sim x^{2/3}$. Biologically, the growth of organisms (populations of cells) has long been known to scale like $f(x)\sim x^{k}$ with $k\simeq 3/4$ [27], which can be understood in terms of hydrodynamic constraints on vascular and pulmonary networks. For reasons that are not currently understood, a similar pattern of growth appears to recur at the level of ecological communities [28, 17]. In a different direction, it has been recently suggested that predator-prey interactions can be modelled with a square root law, i.e. $g(x)\sim h(x)\sim x^{1/2}$ [29, 30].
@@
an elementary calculation shows that stability of the homogeneous equilibrium $x_{i}^{*}=x^{*}$ requires $\psi(x^{*})g(x^{*})h^{\prime}(x^{*})>0$, or $\psi(x^{*})>0$ if we assume that $g$ and $h$ are positive, increasing functions. This stability condition involves the relative strength of diagonal and off-diagonal interactions ($\mu_{s}-\mu$), but also on the relative growth rate of the production and response functions near the equilibrium ($f^{\prime}(x^{*})/f(x^{*})-g^{\prime}(x^{*})/g(x^{*})$).
@@
Random interactions: stability condition.—We now consider the case of random interactions. Specifically, we assume that interaction coefficients $A_{ij}$ are drawn independently from a distribution with mean $\mu$ and standard deviation $\sigma$; We assume the diagonal elements $A_{ii}$ have a mean value $\mu_{s}$ (possibly different from $\mu$) but the same standard deviation $\sigma$.
@@
$J_{ii}^{*}$ $=f^{\prime}(x_{i}^{*})-g^{\prime}(x_{i}^{*})f(x_{i}^{*})/g(x_{i}^{*})-A_{ii}g(x_{i}^{*})h^{\prime}(x_{j}^{*})\ ,$ (5)
@@
$M_{i}=-g(x_{i}^{*})h^{\prime}(x_{i}^{*})\psi(x_{i}^{*}),\quad L=g(x_{i}^{*}),\quad R=h^{\prime}(x_{i}^{*}),$ (7)
where $\psi(x)$ is the function defined in (2).
Figure 1: The equilibrium distribution with power-law self- and cross-interactions is accurately reproduced by Eq. (12) and its Gaussian approximation. For simulations (in green) we used $\alpha=1$, $\beta=3/2$, $\gamma=1$, $N=50$, $\mu=\mu_{s}=10^{-2}$, $\sigma=5\cdot 10^{-3}$.
@@
In the GLV model, $\psi(x)=\mu_{s}-\mu$ (independently of $x$) and we recover the classical condition $\sigma\sqrt{N}<(\mu_{s}-\mu)$  [^1] In fact, the same condition holds when $f$ and $g$ are power laws with the same exponent.. In general, however, the dependence on equilibrium values $x_{i}^{*}$ (which in turn depend on $N$) does not cancel out, and one must gain information about the distribution of equilibrium values $P(x_{i}^{*})$ to assess the stability condition (8).
@@
Figure 2: For a system in the “anti-May” phase, an increase in $N$, corresponding to moving along $\sqrt{N}$ in the $(\sigma\sqrt{N},\mu N)$ plane (green dashed line), will be stabilizing rather than destabilizing. Here we show results from the numerical resolution of the dynamical system (1). Stability is defined as full stable coexistence. The red line is computed obtained using DMFT and the generalized stability condition $N\langle(\sigma/\psi)^{2}\rangle<1$. Parameters values are $\alpha=1$, $\beta=3/2$, $\gamma=1$, $N=50$ and $10$ replicates for the simulations. The green line was plotted for $\mu=0.1$ and $\sigma=0.75$.
Following e.g. Ref. [16], we can derive from the evolution equation (1) a dynamical mean field theory (DMFT) which describes the ensemble of $N\gg 1$ variables by means of a single, self-consistent representative stochastic trajectory
$\dot{x}=x^{\alpha}-A_{s}x^{\beta+\gamma}-x^{\beta}\big{(}\mu N\langle x^{\gamma}\rangle+\sigma\sqrt{N}\eta\big{)}\,,$ (9)
where $A_{s}$ is a random variable with the statistics of $A_{ii}$ and $\eta$ is a Gaussian noise with zero mean and correlation $\langle x^{\gamma}(t)x^{\gamma}(s)\rangle$.
Figure 3: The “May” and “anti-May” phases derived in the case of uniform interactions ($\sigma=0$) are robust with respect to random interactions ($\sigma>0$). Here we simulated system (1) in the power law case for a fixed value of $\mu=\mu_{s}=10^{-2}$, $\gamma=1$, $N=100$ and $100$ replicates, and we evaluated $\sigma_{c}$, defined here as the value above which full stable coexistence is less probable than 50%, at varying $\alpha$ and $\beta$. No amount of heterogeneity is allowed in the bottom right triangle because we have $\mu=\mu_{s}$, while an increasing value of $\sigma_{c}$ is found as $\beta$ becomes larger than $\alpha$. We stress that the existence of a finite $\sigma_{c}\sqrt{N}$ above which instability is triggered also in the upper triangle, does not imply that increasing $N$ will destabilize the system, because $\mu N$ would also change.
At equilibrium, we can write
$x_{*}^{\alpha-\beta}-A_{s}x_{*}^{\gamma}=\big{(}\mu N\langle x_{*}^{\gamma}\rangle+\sigma\sqrt{N\langle x_{*}^{2\gamma}\rangle}\xi\big{)}\,,$ (10)
where $\xi$ is now a standard normal random variable. The equation above can be solved for $x_{*}$ for specific values of $(\alpha-\beta)/\gamma$, or if $A_{s}=0$. (In the following we show that neglecting $A_{s}$ even when it is not zero does not affect noticeably the quality of the approximation.) In this case the stationary solution is given by
$x_{*}=\left(\mu N\langle x_{*}^{\gamma}\rangle+\sigma\sqrt{N\langle x_{*}^{2\gamma}\rangle}\xi\right)^{1/(\alpha-\beta)}\,.$ (11)
@@
where the expectation values must be computed self-consistently.
@@
Equipped with the equilibrium distribution $P(x^{*})$, we can compute, for given $N$ and $\mu$, the maximal heterogeneity $\sigma_{c}$ compatible with the linear stability condition $N\langle(\sigma/\psi(x^{*}))^{2}\rangle<1$.
Let us focus on the choice $\alpha=1$, $\beta=3/2$ and $\gamma=1$, with $A_{s}=0$. In this case the first and second moments formally diverge. This is not truly problematic because we are dealing with large, but finite populations. Following [33, 17] and assuming $\sigma\sqrt{N\langle x_{*}^{2}\rangle}\ll\mu N\langle x_{*}\rangle$, we can expand the solution (11) to first order in $\sigma\sqrt{N\langle x_{*}^{2}\rangle}$, and approximate (12) with a Gaussian distribution with moments $\langle x_{*}\rangle=(\mu N)^{-2/3}$ and $\langle x_{*}^{2}\rangle=(\mu N)^{-4/3}/(1-4(\mu N)^{-2}\sigma^{2}N)$.
The result is plotted against simulations in Fig. 1. We show both the exact solution (12) (where integration for the moments has been performed up to a maximum value above which no equilibrium value should be found) and the Gaussian approximation. Simulations correspond to $A_{s}=\mu$, showing that neglecting the $A_{s}$ term in analytical expressions does not introduce a significant error.
@@
The properties of large random dynamical systems is often portrayed in the $(\sigma\sqrt{N},\mu N)$ plane [25]. Everything else being fixed, an increase of the number of degrees of freedom $N$ moves the system along a square-root trajectory in that plane. In the GLV model, the boundary between the stable and unstable parameter regions is the horizontal line $\sigma_{c}\sqrt{N}=\mu_{s}$ [25], hence stability is never possible at large $N$.
@@
In Fig. 2 we plot the stability condition $N\langle(\sigma/\psi(x^{*}))^{2}\rangle<1$ for $\alpha=1$, $\beta=3/2$ derived from (12) (solid line) together with results from simulations (shading). Here, because the boundary between the stable and unstable phases is a straight line with positive slope, increasing $N$ will eventually bring the system in the stable region. This behavior corresponds to the “anti-May” phase defined previously.
@@
Finally, Fig. 3 shows results of simulations for $\sigma_{c}\sqrt{N}$ in the $(\alpha,\beta)$ plane, at fixed $\mu N$ and $\gamma=1$. (Numerically, we define $\sigma_{c}\sqrt{N}$ as the value of $\sigma$ above which full stable coexistence has a probability lower than $50\%$.) In these simulations we use $\mu_{s}=\mu$, which would never be stable in the GLV model. Fig. 3 illustrates the transition between a “May” phase for $\alpha\geq\beta$ and an “anti-May” phase for $\alpha<\beta$, showing that our analytical results are robust to heterogeneous interactions.
@@
Discussion.—The relationship between complexity and stability in high-dimensional dynamical systems has been a longstanding puzzle, in ecology and in other fields. In this letter, we have showed that the condition $\sigma\sqrt{N}<\mu_{s}-\mu$ (popularized by May’s slogan “complexity begets instability”) does not provide a complete picture of the relationship between complexity and stability. In particular, we have seen that the generalized Lotka-Volterra model, often cited in support of May’s general prediction, exemplifies a special cancellation in the more general stability condition $N\langle(\sigma/\psi)^{2}\rangle<1$. In models where self-interactions grow faster than cross-interactions, the opposite behavior is observed: stability becomes more likely with increasing diversity $N$.

### v2

# Complexity-stability relationships in disordered dynamical systems

@@ Abstract
Robert May famously used random matrix theory to predict that large, complex systems cannot admit stable fixed points. However, this general conclusion is not always supported by empirical observation: from cells to biomes, biological systems are large, complex and, often, stable. In this paper, we revisit May’s argument in light of recent developments in both ecology and random matrix theory. Using a non-linear generalization of the competitive Lotka-Volterra model, we show that there are, in fact, two kinds of complexity-stability relationships in disordered dynamical systems: if self-interactions grow faster with density than cross-interactions, complexity is destabilizing; but if cross-interactions grow faster than self-interactions, complexity is stabilizing.
@@
Introduction.—Few mathematical arguments have influenced biological thinking like May’s prediction that large, complex ecosystems cannot be stable [1]. It is, on the face of it, a perplexing conclusion. On the one hand, May’s mathematical argument is simple and seemingly model-free, suggesting universal applicability, also beyond biology [2, 3]. On the other hand, it is clear that at least some large, complex systems are stable—else which regularities would biology be studying in the first place? In fact, empirical observation suggests the opposite relationship between complexity and stability: species-rich, strongly-coupled communities such as rainforests tend to be stable over time, while sparser ones, for instance arctic communities, often exhibit large fluctuations, extinctions, and invasions [4, 5, 6]. The tension between May’s theoretical argument and observation is at the center of the longstanding “diversity-stability debate” in ecology [7, 8, 9].
@@
May’s argument can be summarized as follows [1]. Consider a system with $N$ populations $x_{i}$, characterized by an equilibrium point $\mathbf{x}^{*}$. Near that equilibrium $\mathbf{x}=\mathbf{x}^{*}+\delta\mathbf{x}$, the dynamics of the system is described by linear equations $d(\delta\mathbf{x})/dt=A(\delta\mathbf{x})$, and the stability of these equations requires that all eigenvalues of $A$ have negative real part. But if $A$ can be represented as $A=B-I$, where $B$ consists of random, independent interactions (with zero mean and variance $\sigma^{2}$), and $-I$ corresponds to stabilizing self-interactions on some natural timescale, the circular law of random matrix theory implies that all eigenvalues of $A$ will have negative real part only if $\sigma^{2}N<1$. This condition places a sharp constraint on both diversity $N$ and interaction strength $\sigma$, often referred to as “complexity begets instability”. (This argument generalizes to $\langle B_{ij}\rangle\neq 0$, incomplete connectivity, or correlated interactions [10].)
@@
Many authors have sought to ease the tension between May’s prediction and empirical observation by invoking effects not captured by dynamical systems with random coefficients [7, 11, 12, 13, 14, 15]. In this Letter, we use recent results in the physics of disordered systems [16, 17] to show that May’s argument itself is incomplete: in random dynamical systems, stability does not necessarily decrease with dimensionality and interaction strength—the opposite behavior is also possible, without the need for special or additional structure. For an in-depth discussion of the ecological implications of our result, see our recent paper [9].
@@
Here $f(x_{i})$ represents the self-dynamics of a population $i$ (growth and self-regulation), while $g(x_{i})$ and $h(x_{j})$ capture the cross-interaction of $i$ with other populations. That interaction is weighted by a coefficient $A_{ij}>0$, such that it implies a negative effect of $j$ on the growth of $i$. Non-zero diagonal elements $A_{ii}$ accounts for additional self-interactions with the same functional form as cross-interactions. We refer to $f$ as the “production function” and $g$ as the “coupling function”.
This general setting has been used to study universality in network dynamics [18] and to construct minimal models of complex dynamics [19], with applications ranging from biological [20, 21] to social [22, 23, 24] systems. In particular, the classic generalized Lotka-Volterra (GLV) competition model studied by Bunin and collaborators [25, 26] corresponds to $f$, $g$, $h$ all linear. However, it is natural both physically and biologically to consider more general functions, including power laws $f(x)\sim x^{\alpha}$, $g(x)\sim x^{\beta}$, $h(x)\sim x^{\gamma}$. From a physical perspective, we can imagine populations $x_{i}$ forming three-dimensional clusters whose growth is limited to their two-dimensional surface, leading to a production function $f(x)\sim x^{2/3}$. Biologically, at the individual organism level (populations of cells), growth has long been known to scale like $f(x)\sim x^{k}$ with $k\simeq 3/4$ [27], which can be understood in terms of hydrodynamic constraints on vascular and pulmonary networks. For reasons that are not currently understood, a similar pattern of growth appears to recur at the level of ecological communities [28, 9]. In a different direction, it has been recently suggested that predator-prey interactions can be modelled with a square root law, i.e. $g(x)\sim h(x)\sim x^{1/2}$ [29, 30].
@@
an elementary calculation shows that stability of the homogeneous equilibrium $x_{i}^{*}=x^{*}$ requires $\psi(x^{*})g(x^{*})h^{\prime}(x^{*})>0$, or $\psi(x^{*})>0$ if we assume that $g$ and $h$ are positive, increasing functions. This stability condition involves the relative strength of diagonal and off-diagonal interactions ($\mu_{s}-\mu$), but also on the relative growth rate of the production and coupling functions near the equilibrium ($f^{\prime}(x^{*})/f(x^{*})-g^{\prime}(x^{*})/g(x^{*})$).
@@
Random interactions: stability condition.—We now consider the case of random interactions. Specifically, we assume that interaction coefficients $A_{ij}$ are drawn independently from a distribution with mean $\mu$ and standard deviation $\sigma$. We assume the diagonal elements $A_{ii}$ have a mean value $\mu_{s}$ (possibly different from $\mu$) and standard deviation $\sigma$.
@@
$J_{ii}^{*}$ $=f^{\prime}(x_{i}^{*})-g^{\prime}(x_{i}^{*})f(x_{i}^{*})/g(x_{i}^{*})-A_{ii}g(x_{i}^{*})h^{\prime}(x_{i}^{*})\ ,$ (5)
@@
$M_{ij}=-g(x_{i}^{*})$ $h^{\prime}(x_{i}^{*})\psi(x_{i}^{*})\delta_{ij},$
$L_{ij}=g(x_{i}^{*})\delta_{ij},$ $\quad R_{ij}=h^{\prime}(x_{i}^{*})\delta_{ij},$ (7)
where $\delta_{ij}$ is the Kronecker delta and $\psi(x)$ is the function defined in (2).
Figure 1: The equilibrium distribution for the case of power-law self- and cross-interactions is accurately reproduced by Eq. (11). For simulations (in green) we used $\alpha=1$, $\beta=3/2$, $\gamma=1$, $N=100$, $\mu=\mu_{s}=\sigma=10^{-2}$.
@@
In the GLV model, $\psi(x)=\mu_{s}-\mu$ (independently of $x$) and we recover the classical condition $\sigma\sqrt{N}<\mu_{s}-\mu$  [^1] In fact, the same condition holds when $f$ and $g$ are power laws with the same exponent.. In general, however, the dependence on equilibrium values $x_{i}^{*}$ (which in turn depends on $N$) does not cancel out, and one must gain information about the distribution of equilibrium values $P(x_{i}^{*})$ to assess the stability condition (8).
@@
Figure 2: For a system in the “anti-May” phase, an increase in $N$, corresponding to moving along $\sqrt{N}$ in the $(\sigma\sqrt{N},\mu N)$ plane (green dashed line), will be stabilizing rather than destabilizing. Here we show results from the numerical resolution of the dynamical system (1). Stability is defined as full stable coexistence. The red line is computed using the cavity solution and the generalized stability condition $N\langle(\sigma/\psi)^{2}\rangle<1$. Parameters values are $\alpha=1$, $\beta=3/2$, $\gamma=1$, $N=50$ and $10$ replicates for the simulations. The green line is plotted for $\mu=0.1$ and $\sigma=0.75$.
Following e.g. Ref. [17], we can derive from Eq. (1) a cavity solution that describes the ensemble of the stationary solutions of the system by means of a single representative random variable $x_{*}$
$0=x_{*}^{\alpha}-A_{s}x_{*}^{\beta+\gamma}-x_{*}^{\beta}\big{(}\mu N\langle x_{*}^{\gamma}\rangle+\sigma\sqrt{N}\sqrt{\langle x_{*}^{2\gamma}\rangle}\xi\big{)}\,,$ (9)
where $A_{s}$ is a random variable with the statistics of $A_{ii}$ and $\xi$ is a standard normal random variable.
Figure 3: The “May” and “anti-May” phases derived in the case of uniform interactions ($\sigma=0$) are robust with respect to random interactions ($\sigma>0$). Here we simulated system (1) in the power law case for a fixed value of $\mu=\mu_{s}=10^{-2}$, $\gamma=1$, $N=100$ and $100$ replicates, and we evaluated $\sigma_{c}$, defined here as the value above which full stable coexistence is less probable than 50%, at varying $\alpha$ and $\beta$. No amount of heterogeneity is allowed in the bottom right triangle because we have set $\mu=\mu_{s}$, while an increasing value of $\sigma_{c}$ is found as $\beta$ becomes larger than $\alpha$. We stress that the existence of a finite $\sigma_{c}\sqrt{N}$ above which instability is triggered also in the upper triangle, does not imply that increasing $N$, at fixed $\mu$ and $\sigma$, will destabilize the system, because $\mu N$ would also change.
We are interested in the case in which the number of degrees of freedom $N$ is the only parameter changing in the system and the strength and heterogeneity of the interactions are fixed [33, 34]. Compared to the case where $\mu$ and $\sigma$ scale with $N$ (considered in [25] and follow-up works), our choice reflects a strong-interaction regime; it has been employed, for example, to study ecological scenarios with large differences in species abundance and chaotic turnover [35]. As we show below, the cavity approximation does a good job at describing the effects of heterogeneity in the interactions as long as $\sigma\sqrt{N}$ is not negligible relative to $\mu N$, at which point we recover the homogeneous case.
The equation above can be solved for $x_{*}$ for specific values of $(\alpha-\beta)/\gamma$, or if $A_{s}=0$. (In the following we show that neglecting $A_{s}$, when it is of the order of $\mu$, does not affect noticeably the quality of the approximation.) In this case the stationary solution is given by
$x_{*}=\left(\mu N\langle x_{*}^{\gamma}\rangle+\sigma\sqrt{N}\sqrt{\langle x_{*}^{2\gamma}\rangle}\xi\right)^{1/(\alpha-\beta)}\,.$ (10)
@@
defined from 00 to $\infty$, and where the expectation values must be computed self-consistently. Equipped with the equilibrium distribution $P(x_{*})$, we can compute, for given $N$ and $\mu$, the maximal heterogeneity $\sigma_{c}$ compatible with the linear stability condition $N\langle(\sigma/\psi(x_{*}))^{2}\rangle<1$.
[…]
@@
The properties of large random dynamical systems are often portrayed in the $(\sigma\sqrt{N},\mu N)$ plane [25]. If we keep the mean $\mu$ and standard deviation $\sigma$ of interactions fixed, an increase in the number of degrees of freedom $N$ moves the system along a square-root trajectory in that plane. In the GLV model, the boundary between the stable and unstable parameter regions is the horizontal line $\sigma_{c}\sqrt{N}=\mu_{s}-\mu N/N\approx\mu_{s}$ [25], hence stability is never possible at large $N$.
@@
In Fig. 2 we plot the stability condition $N\langle(\sigma/\psi(x_{*}))^{2}\rangle<1$ for $\alpha=1$, $\beta=3/2$ derived from (11) (solid line) together with results from simulations (shading). Here, because the boundary between the stable and unstable phases is a straight line with positive slope, increasing $N$ will eventually bring the system in the stable region. This behavior corresponds to the “anti-May” phase defined previously.
@@
Fig. 3 shows results of simulations for $\sigma_{c}\sqrt{N}$ in the $(\alpha,\beta)$ plane, at fixed $\mu N$ and $\gamma=1$. (Numerically, we define $\sigma_{c}$ as the value of $\sigma$ above which full stable coexistence has a probability lower than $50\%$.) In these simulations, we use $\mu_{s}=\mu$, which would never be stable in the GLV model. Fig. 3 illustrates the transition between a “May” phase for $\alpha\geq\beta$ and an “anti-May” phase for $\alpha<\beta$, showing that our analytical results are robust to heterogeneous interactions.
Finally, we check the robustness of our findings in the scenario in which the exponents are not the same for every variable $x_{i}$. As a test, we can sample exponents from different Gaussian distributions: $\alpha_{i}\sim\mathcal{N}(\alpha,\sigma_{e})$, $\beta_{i}\sim\mathcal{N}(\beta,\sigma_{e})$ and $\gamma_{i}\sim\mathcal{N}(\gamma,\sigma_{e})$, respectively with mean $\alpha$, $\beta$ and $\gamma$, and, for simplicity, all with the same standard deviation $\sigma_{e}$. Simulations show that our results hold as long as $\sigma_{e}$ is small enough, in particular, we observe a loss of stability when $\min_{i}\beta_{i}>\max_{i}\alpha_{i}$ (Supplemental Material).
@@
Discussion.—The relationship between complexity and stability in high-dimensional dynamical systems has been a longstanding puzzle, in ecology and other fields. In this Letter, we have shown that the condition $\sigma\sqrt{N}<\mu_{s}$ does not provide a complete picture of the relationship between complexity and stability. In particular, we have seen that the competitive Lotka-Volterra model, often cited in support of May’s general prediction, exemplifies a special cancellation in the more general stability condition $N\langle(\sigma/\psi)^{2}\rangle<1$. In models where self-interactions grow slower than cross-interactions [9, 37], the opposite behavior is observed: stability becomes more likely with increasing diversity $N$.
A modification of our model (1) of general ecological relevance describing the case in which per capita growth rates depend on a linear combination of the densities is $\dot{x_{i}}/x_{i}=1-F(\sum_{j}A_{ij}x_{j})$, where the function $F$ characterizes the effects of competition. General models of this kind are used, for example, to study the growth of a clonal population in the presence of a limiting factor [38]. All the models in this class are destabilized by diversity. (The Jacobian matrix at equilibrium is $J_{ij}^{*}=-A_{ij}x_{i}^{*}F^{\prime}(\sum_{j}A_{ij}x_{j}^{*})$, leading to the same simplification of the stability condition (8) as in the GLV model.)